OpenStudy (anonymous):

Perform the Indicated operations. If the matrix does not exist, write impossible.

OpenStudy (anonymous):

let me just insert the matrix...

OpenStudy (turingtest):

c

OpenStudy (anonymous):

\[2\left[\begin{matrix}-1 \\ 5 \\ -6\end{matrix}\right]+4\left[\begin{matrix} -3x \\ 2 \\ x\end{matrix}\right]-3\left[\begin{matrix}x-2 \\ 3 \\ 1\end{matrix}\right]\]

OpenStudy (anonymous):

⎡ 4-15x ⎤ | 9 | ⎣-15+4x ⎦

OpenStudy (anonymous):

how did u get that?

OpenStudy (turingtest):

the top row: 2(-1)+4(-3x)-3(x-2)=-2-12x-3x-6=4-15x middle row: 2(5)+4(2)-3(3)=10+8-9=9 bottom row: 2(-6)+4(x)-3(1)=-12+4x-3=15+4x i.e.) multiply every element of each matrix by the scalar in front of it, then add straight across like regular algebra.

OpenStudy (anonymous):

wouldnt -3(x-2) = -3x+6, not minus?

OpenStudy (anonymous):

never mind thank you :D